What is the sum of -13 + 5 + (-7) ?
Is it -15 ?

Answer by lenny460(1073)

What is the sum of -13 + 5 + (-7) ?
Is it -15 ?
-13 + 5 = -8
-8 + (-7) = -15


Answer:
-15